Only two gallium isotopes are naturally occurring, 69Ga (mass = 68.9256 amu) and 71Ga (mass = 70.9247 amu). The atomic mass of gallium is 69.72. What percent of the gallium atoms (% abundance) is the heavier isotope?

[1] 36 %

[2] 38 %

[3] 40 %

[4] 59 %

[5] 61 % 


1
Expert's answer
2022-03-19T08:52:40-0400

We can write such system of equations

x+y=1

68.9256x+70.9247y=69.72


x - part of the 69Ga

y- part of the 71Ga


From first equalation we havex=1-y

68.9256(1-y)+70.9247y=69.72

1.9991 y=0.7944

y=0.4=40%

Answer: [3] 40%